22 XTMale eu) (013) (0%) whn Male iady variable Male1 o tme male resianestimates Calcula e ussion SER
Answer #1

(a) We may see that Female, 1-Male, or Male, 1-Female, . Putting this into the regression llage, 12.52+2.12 * Male, , we have llage, = 12.52+2.12* (1-Fernale, or llage, = 12.52+2.12-2.12 * Female, or 11 age, 1 4.64-2. 1 2 * Female . The residual standard error (SER) and the r-square would also be the same as before, since the residual sum of square would be same as before.

The answer would be hence 11.age, 14.64-2. Î2 * Femaleï ,

R-=( and SER=. .